Can someone help me on this problem?
I need to find the GCF of 6, 18 and 30.
Thank you.

Answer by funmath(2925)

6 is the biggest number that divides evenly into all three numbers.
6/6=1
18/6=3
30/6=5
Another way to do it is to find the one number that is a factor in all three numbers:
6=6*1
18=6*3
30=6*5
Another way is to find the prime factors using a factor tree:
6=2*3
18=2*3*3
30=2*3*5
They all have 2*3 in common. 2*3=6